OpenStudy (anonymous):

Find the side length of a square with the given area. 121 in² A. 11 in. B. 60.5 in. C. 1331 in. D. 14,641 in

jimthompson5910 (jim_thompson5910):

Solve for s \[\Large A = s^2\] \[\Large 121 = s^2\] \[\Large s = ???\]

OpenStudy (anonymous):

it's c right

jimthompson5910 (jim_thompson5910):

how are you getting 1331?

OpenStudy (anonymous):

I meant to say B